OpenStudy (amistre64):

so, before i forget with the question on the exam was .... What is the equilibrium temp if you mix 500 grams of 350K water, with 350 grams of 300K water? How would you solve it, and was 329.41K a good answer?

OpenStudy (anonymous):

I would solve it by using the fact that the heat lost by the 350K water=the heat gained by the 300K water. i.e. \(0.5*4186*(350-\theta)=0.35*4186*(\theta-300)\) That gave me the exact same answer as you. So everything's good :)

OpenStudy (amistre64):

yay!! that makes me happier about my answer :)